# Planning of Off-Grid Renewable Power to Ammonia Systems with Heterogeneous Flexibility: A Multistakeholder Equilibrium Perspective

### v1

@@ Abstract
Off-grid renewable power to ammonia (ReP2A) systems present a promising pathway toward carbon neutrality in both the energy and chemical industries. However, due to chemical safety requirements, the limited flexibility of ammonia synthesis poses a challenge when attempting to align with the variable hydrogen flow produced from renewable power. This necessitates the optimal sizing of equipment capacity for effective and coordinated production across the system. Additionally, an ReP2A system may involve multiple stakeholders with varying degrees of operational flexibility, complicating the planning problem. This paper first examines the multistakeholder sizing equilibrium (MSSE) of the ReP2A system. First, we propose an MSSE model that accounts for individual planning decisions and the competing economic interests of the stakeholders of power generation, hydrogen production, and ammonia synthesis. We then construct an equivalent optimization problem based on Karush–Kuhn–Tucker (KKT) conditions to determine the equilibrium. Following this, we decompose the problem in the temporal dimension and solve it via multicut generalized Benders decomposition (GBD) to address long-term balancing issues. Case studies based on a realistic project reveal that the equilibrium does not naturally balance the interests of all stakeholders due to their heterogeneous characteristics. Our findings suggest that benefit transfer agreements ensure mutual benefits and the successful implementation of ReP2A projects.
Index Terms:
renewable power to ammonia, sizing, investment equilibrium, multi-stakeholder interests, green hydrogen
@@
$mn$
Index for the pipelines between nodes $m$ and $n$
@@
Active/reactive power flows on branch $ij$
@@
Square of current on branch $ij$
@@
Square of the voltage amplitude at bus $j$
@@
Active and reactive power injections at bus $j$
@@
Average hydrogen flow of pipeline $mn$
@@
Hydrogen inflow/outflow of pipeline $mn$
@@
Linepack storage of pipeline $mn$
@@
$\underline{\overline{W}}^{\text{rg,wt/pv/bes/vc}}$
WT/PV/BES/VC capacity limits of the RG
$\underline{\overline{W}}^{\text{hp,bes/ae/hst}}$
BES/AE/HST capacity limits of the HP
$\underline{\overline{W}}^{\text{as,hst/asy/ast}}$
HST/ASY/AST capacity limits of the AS
@@
¯$\overline{\ell}_{ij}$
Current limit of branch $ij$
@@
Weymouth constants of piepline $mn$
@@ I-A Background and Motivation
Renewable power to ammonia (ReP2A) offers a promising pathway for the large-scale utilization of renewable energy sources (RESs) and carbon emission reduction in the power and chemical industries [1, 2]. ReP2A projects have been implemented in many countries, including China [3], Denmark, and Australia [4]. Off-grid systems, in particular, have shown greater potential due to policy support [5] and flexibility in planning without grid integration constraints [6].
@@
However, involving multiple investors [7] presents additional challenges in some projects. For example, a renewable power-to-hydrogen (P2H) plant might be invested in by two stakeholders (renewable promoters and fertilizer producers) [9]. In some cases, investments in ammonia plants do not include facilities for hydrogen production and renewable power generation [10], leading to the ASY investor being a separate stakeholder. Consequently, the ReP2A system may involve three stakeholders: renewable power generation, hydrogen production, and ammonia synthesis (denoted RG, HP, and AS, respectively). These stakeholders often have conflicting interests, and their planning and operations influence each other.
@@ I-B Literature Review
However, the aforementioned studies consider the ReP2A system as planned and operated by a single entity. In the presence of multiple stakeholders and competition in transactions, unified planning and operation may not be feasible. Research on multistakeholder equilibrium [20, 21, 22] have addressed pricing and trading in day-ahead markets for power systems [20], microgrids [21], and multienergy systems [22]. However, the ReP2A system differs significantly due to its heterogeneous flexibility among stakeholders with staged electricity–hydrogen–ammonia production processes and BSs. Additionally, existing research has focused primarily on operational aspects, with limited analysis of planning issues. Overall, there is currently insufficient research on multistakeholder sizing equilibrium (MSSE) in ReP2A systems leaving a gap in investment equilibrium in the green ammonia industry.
@@ II Structure of ReP2A Systems
In the complete green ammonia production process, maintaining energy and mass balance is crucial. The RG, HP, and AS exhibit distinct flexibility due to their inherent physical characteristics. To accommodate the heterogeneous flexibility, multi-type buffer systems (BSs) are implemented, with staged BES, HST, and AST designated for different timescales.
@@
4) The electricity transmission network is owned by the RG, and the hydrogen delivery network is owned by the HP. The network distances are determined by the geographical locations of the RG, HP, and AS.
@@ III-A Modeling the Whole Process of ReP2A
The entire ReP2A process involves the flow of energy and mass related to electricity, hydrogen, and ammonia. The detailed physical models and planning and operational constraints of the corresponding stakeholders are as follows, where 1)–3) pertains to the RG, 5)–8) pertains to the HP, 10)–12) pertains to the AS, and 4) and 9) are the constraints coupling multiple stakeholders.
@@ III-A1 RG Planning Capacities
$\big{[}\underline{W}^{\text{rg,bes}},\underline{W}^{\text{rg,vc}}\big{]}\leq\big{[}W^{\text{rg,bes}},W^{\text{rg,vc}}\big{]}\leq\big{[}\overline{W}^{\text{rg,bes}},\overline{W}^{\text{rg,vc}}\big{]}.$ (1)
@@ III-A3 Electrical Power Flow
The electrical network is described via the DistFlow model [24], as shown in (7)–(11). The power injection at each bus $i$ is determined by (12)–(16).
@@ III-A4 Electricity Transactions
$\begin{cases}P_{t}^{\text{rg,sell,hp}}-P_{t}^{\text{hp,buy,rg}}=0:(\rho_{t}^{\text{rg-hp,e,rg}},\rho_{t}^{\text{rg-hp,e,hp}},\rho_{t}^{\text{rg-hp,e,as}}),\\ P_{t}^{\text{rg,sell,as}}-P_{t}^{\text{as,buy,rg}}=0:(\rho_{t}^{\text{rg-as,e,rg}},\rho_{t}^{\text{rg-as,e,hp}},\rho_{t}^{\text{rg-as,e,as}}),\end{cases}$ (17)
@@ III-A7 HST Operation
The HST invested in by the HP stakeholder is used to accommodate the varying hydrogen flow and maximize the benefit from the hydrogen transactions with the AS stakeholder. The state of the HST satisfies (23)–(25), with hydrogen charging and release constrained by (26).
@@
$[0,0]\leq[f_{t}^{\text{hp,hst,in}},f_{t}^{\text{hp,hst,out}}]\leq[0.5W^{\text{hp,hst}},0.5W^{\text{hp,hst}}].$ (26)
@@ III-A8 Hydrogen Delivery Network
$f_{t}^{\text{hp,pro}}+f_{t}^{\text{hp,hst,out}}-f_{t}^{\text{hp,hst,in}}+F_{mn,t}^{\text{out}}-F_{mn,t}^{\text{in}}=f_{t}^{\text{hp,sell,as}}.$ (33)
@@ III-B Multi-Stakeholder Sizing Model
$C_{\text{ope}}^{\text{rg}}=$ $\frac{8760}{T}\textstyle\sum_{t=1}^{T}(\sigma^{\text{deg}}P_{t}^{\text{rg,bes,d}}-P_{t}^{\text{rg,sell,hp}}\rho_{t}^{\text{rg-hp,e,rg}})\Delta t$
@@
where $\text{CRF}(r,y)={r(1+r)^{y}}/{(\left(1+r\right)^{y}-1)}$ is the capital recovery factor; $r=8\%$ and $y$ are the discount rate and lifetime; $C^{\text{rg,line}}$ is the transmission line cost; and ${\eta}_{\text{O\&M}}=2\%$ is the annual operation and maintenance (O&M) cost factor.
@@
$C_{\text{inv}}^{\text{as}}=$ $\text{CRF}(r,y)(c^{\text{as}}W^{\text{as,asy}}+c^{\text{ast}}W^{\text{as,ast}}+c^{\text{hst}}W^{\text{as,hst}}),$ (53)
$C_{\text{ope}}^{\text{as}}=$ $\frac{8760}{T}\sum\nolimits_{t=1}^{T}\big{[}(f_{t}^{\text{as,buy,hp}}\rho_{t}^{\text{hp-as,h,as}}+P_{t}^{\text{as,buy,rg}}\rho_{t}^{\text{rg-as,e,as}})\Delta t$
@@
Given the above, we summarize the multistakeholder planning problem in Table I. For simplicity, we use $(\bm{x}^{k})_{k\in\{\text{rg,hp,as}\}}$ and $(\bm{g}^{k})_{k\in\{\text{rg,hp,as}\}}$ to denote all independent decision variables and constraints of the RG, HP, and AS.
Considering the individual rationality of each stakeholder and competition in the transactions, we frame the multi-stakeholder sizing problem as a non-cooperative game and propose an MSSE model for the ReP2A system, with the game structure depicted in Fig. 2. According to (46)-(54), there are conflicts of interest among RG, HP, and AS, particularly in terms of the revenues and costs related to electricity and hydrogen trading, which will be discussed in Section 10. Moreover, the investment costs of each stakeholder are not balanced due to heterogeneous physical characteristics in the staged processes, and we discuss this later in Sections V-C and V-D.
@@ IV-A Direct Approach Based on KKT Conditions
The equilibrium of noncooperative games is usually solved via fixed-point iteration [26] and direct approaches [27, 22], including those based on Karush–Kuhn–Tucker (KKT) conditions [27] and strong duality [22]. However, the fixed-point iteration approach cannot ensure convergence, and strong duality introduces bilinear terms, making it challenging to solve. Thus, we attempt to solve the MSSE model via KKT conditions.
@@
$\nabla_{\bm{x}^{k}}C^{k}+(\bm{\lambda}^{k})^{\text{T}}\nabla_{\bm{x}^{k}}\bm{g}^{k}+(\bm{\rho}^{\text{rg-hp,e},k})^{\text{T}}\nabla_{\bm{x}^{k}}(P^{\text{rg,sell,hp}}-P^{\text{hp,buy,rg}})$
$+(\bm{\rho}^{\text{hp-as,h},k})^{\text{T}}\nabla_{\bm{x}^{k}}(f^{\text{hp,sell,as}}-f^{\text{as,buy,hp}})=0,$ (55a)
$(\bm{\lambda}^{k})^{\text{T}}\bm{g}^{k}=0,~{}\bm{\lambda}^{k}\geq\bm{0},~{}\bm{g}^{k}\leq\bm{0},$ (55b)
[…]
where $\bm{\lambda}^{k}$ is the dual variable of $\bm{g}^{k}$.

### v2

@@ Abstract
Off-grid renewable power to ammonia (ReP2A) systems present a promising pathway toward carbon neutrality in both the energy and chemical industries. However, due to chemical safety requirements, the limited flexibility of ammonia synthesis poses a challenge when attempting to align with the variable hydrogen flow produced from renewable power. This necessitates the optimal sizing of equipment capacity for effective and coordinated production across the system. Additionally, an ReP2A system may involve multiple stakeholders with varying degrees of operational flexibility, complicating the planning problem. This paper first examines the multistakeholder sizing equilibrium (MSSE) of the ReP2A system. First, we propose an MSSE model that accounts for individual planning decisions and the competing economic interests of the stakeholders of power generation, hydrogen production, and ammonia synthesis. We then construct an equivalent optimization problem based on Karush–Kuhn–Tucker (KKT) conditions to determine the equilibrium. Following this, we decompose the problem in the temporal dimension and solve it via multicut generalized Benders decomposition (GBD) to address long-term balancing issues. Case studies based on a realistic project reveal that the equilibrium does not naturally balance the interests of all stakeholders due to their heterogeneous characteristics. Our findings suggest that benefit transfer or re-arrangement ensure mutual benefits and the successful implementation of ReP2A projects.
@@
mnmn
Index for the pipeline between nodes $m$ and $n$
@@
Active/reactive power flows on branch ijij
@@
Square of current on branch ijij
@@
Square of the voltage amplitude at bus $i$
@@
Active and reactive power injections at bus $i$
@@
Average hydrogen flow of pipeline mnmn
@@
Hydrogen inflow/outflow of pipeline mnmn
@@
Linepack storage of pipeline mnmn
@@
${\overline{W}}^{\text{rg,wt/pv/bes/vc}}$
Upper capacity limit of WT/PV/BES/VC in the RG
${\overline{W}}^{\text{hp,bes/ae/hst}}$
Upper capacity limit of BES/AE/HST in the HP
${\overline{W}}^{\text{as,hst/asy/ast}}$
Upper capacity limit of HST/ASY/AST in the AS
$\underline{W}^{\text{rg,wt/pv/bes/vc}}$
Lower capacity limit of WT/PV/BES/VC in the RG
$\underline{W}^{\text{hp,bes/ae/hst}}$
Lower capacity limit of BES/AE/HST in the HP
$\underline{W}^{\text{as,hst/asy/ast}}$
Lower capacity limit of HST/ASY/AST in the AS
@@
$\overline{\ell}_{ij}$
Current limit of branch ijij
@@
Weymouth constants of pipeline mnmn
@@ I-A Background and Motivation
In recent years, wind and solar power installations have rapidly increased, accompanied by significant renewable energy sources (RESs) curtailment. Power to hydrogen (P2H) offers flexible adjustment capabilities, enabling extensive integration of RESs [1]. According to the International Energy Agency (IEA), global installed electrolyzer capacity could reach 5 GW by 2024 and is expected to reach 230 GW by 2030 [2]. Meanwhile, ammonia, an important global chemical material, is produced in hundreds of millions of tons annually [3]. However, traditional fossil fuel-based “grey ammonia” results in high carbon emissions [3]. Renewable power to ammonia (ReP2A) is a promising pathway for the large-scale utilization of RESs and decarbonization in the power and chemical industries [4, 3, 5]. ReP2A projects have been implemented in many countries, including China [6], Denmark, and Australia [7]. To the best of our knowledge, nearly 25 ReP2A projects were initiated in China in 2023 [8].
ReP2A systems can be classified into grid-connected and off-grid types based on their operational modes [9, 6]. Grid-connected systems are more stable but may face challenges such as failing to meet green certification [10, 11] due to the use of fossil fuel-based hydrogen, as well as high costs of connecting to the grid and operational limitations [12]. Off-grid systems, in particular, have shown greater potential due to policy support [13] and flexibility in planning without grid integration constraints [11].
@@
However, involving multiple investors [14] presents additional challenges in some projects. For example, a renewable power-to-hydrogen (P2H) plant might be invested in by two stakeholders (renewable promoters and fertilizer producers) [16]. In some cases, investments in ammonia plants do not include facilities for hydrogen production and renewable power generation [17], leading to the ASY investor being a separate stakeholder. Moreover, many emerging ReP2A projects have been jointly funded by multiple entities. In the project in Narisong, Inner Mongolia, China [18], the hydrogen plant is funded by China Three Gorges Corporation, while the ammonia plant is invested in by Inner Mongolia Yidong Group Jiuding Chemical Co., Ltd. The project in Ruijin, Jiangxi, China [19] is jointly invested by Datang Corporation Ltd., the local government, and HFG Hydrogen Energy Technology Co., Ltd. Other multistakeholder investment projects also include those in Paradip, India [20], and Dubai, UAE [21], which are not elaborated here. Consequently, the ReP2A system may involve three stakeholders: renewable power generation, hydrogen production, and ammonia synthesis (denoted RG, HP, and AS, respectively). These stakeholders often have conflicting interests, and their planning and operations influence each other.
To guide the development of multistakeholder ReP2A systems, some regions in China have implemented policies [6] that suggest an entity to invest in renewable generation and hydrogen production, while the ammonia plant can be invested in by another entity. However, such projects face challenges like the simultaneous initiation and commission of renewable generation and hydrogen production projects [6], and the need to determine downstream hydrogen application [12]. In regions without policy requirements, renewable generation and hydrogen production can be commissioned separately. Furthermore, recent policies [22] in Inner Mongolia, China, allow the construction of multistakeholder ReP2A projects, and encourage stakeholders to sign long-term contracts. While the authority is guiding the development of ReP2A, there is still room for improvement of policies and measures.
@@ I-B Literature Review
However, the aforementioned studies consider the ReP2A system as planned and operated by a single entity. In the presence of multiple stakeholders and competition in transactions, unified planning and operation may not be feasible. Research on multistakeholder equilibrium [32, 33, 34] have addressed pricing and trading in day-ahead markets for power systems [32], microgrids [33], and multienergy systems [34, 35]. However, the ReP2A system differs significantly due to its heterogeneous flexibility among stakeholders with staged electricity–hydrogen–ammonia production processes and BSs. Additionally, existing research has focused primarily on operational aspects, with limited analysis of planning issues. Overall, there is currently insufficient research on multistakeholder sizing equilibrium (MSSE) in ReP2A systems leaving a gap in investment equilibrium in the green ammonia industry.
@@ II Structure of ReP2A Systems
In the complete green ammonia production process, maintaining energy and mass balance is crucial. The RG, HP, and AS exhibit distinct flexibility due to their inherent physical characteristics. To accommodate the heterogeneous flexibility, multi-type buffer systems (BSs) are implemented, with staged BES, HST, and AST designated for different timescales. Additionally, the BES on the source side uses grid-forming technology [36] to provide frequency response, voltage support [37], and enhance system strength [38] ensuring the stable operation of the off-grid system.
@@
4) The electricity transmission network to HP and AS sides is invested in and owned by the RG, as policy recommendations [22] suggest that renewable energy projects can be supported by accompanying transmission infrastructure, while the hydrogen delivery network is owned by the HP. The network distances are determined by the geographical locations of the RG, HP, and AS.
@@ III-A Modeling the Whole Process of ReP2A
The entire ReP2A process involves the flow of energy and mass related to electricity, hydrogen, and ammonia. The BES is used for energy shifting, maintaining power balance, and supporting hydrogen production, while the HST ensures a stable hydrogen supply for ammonia synthesis, and the AST smooths fluctuations in ammonia production. These systems gradually transform fluctuating renewable energy into stable ammonia production. Additionally, BES in the RG and HST in the HP are used to achieve more profits in transactions. The detailed physical models and planning and operational constraints of the corresponding stakeholders are as follows, where 1)–3) pertains to the RG, 5)–8) pertains to the HP, 10)–12) pertains to the AS, and 4) and 9) are the constraints coupling multiple stakeholders.
@@ III-A1 RG Planning Capacities
$\hskip-1.0pt\big[\underline{W}^{\text{rg,bes}},\underline{W}^{\text{rg,vc}}\big]\leq\big[W^{\text{rg,bes}},W^{\text{rg,vc}}\big]\leq\big[\overline{W}^{\text{rg,bes}},\overline{W}^{\text{rg,vc}}\big].\hskip-2.0pt$ (1)
@@ III-A3 Electrical Power Flow
The ReP2A network security constraints are met through power flow constraints. From the perspective of power flow direction and topology, the ReP2A system resembles a distribution network [9], as it features a radial network without a looped structure. The electrical network is described via a branch flow model called DistFlow [40], as shown in (7)–(11). The power injection at each bus $i$ is determined by (12)–(16). We do not use the DC power flow, which is commonly applied in the transmission network, because the DistFlow model is more accurate and it has manageable computational complexity.
@@ III-A4 Electricity Transactions
$\hskip 0.0pt\begin{cases}P_{t}^{\text{rg,sell,hp}}-P_{t}^{\text{hp,buy,rg}}=0:(\rho_{t}^{\text{rg-hp,e,rg}},\rho_{t}^{\text{rg-hp,e,hp}},\rho_{t}^{\text{rg-hp,e,as}}),\\ P_{t}^{\text{rg,sell,as}}-P_{t}^{\text{as,buy,rg}}=0:(\rho_{t}^{\text{rg-as,e,rg}},\rho_{t}^{\text{rg-as,e,hp}},\rho_{t}^{\text{rg-as,e,as}}),\end{cases}\hskip 0.0pt$ (17)
@@ III-A7 HST Operation
The HST invested in by the HP stakeholder is used to accommodate the varying hydrogen flow and maximize the benefit from the hydrogen transactions with the AS stakeholder. The state of the HST satisfies (23)–(25), with hydrogen charging and release being constrained by (26).
@@
$\bm{0}\leq[f_{t}^{\text{hp,hst,in}},f_{t}^{\text{hp,hst,out}}]\leq[0.5W^{\text{hp,hst}},0.5W^{\text{hp,hst}}].$ (26)
@@ III-A8 Hydrogen Delivery Network
$f_{t}^{\text{hp,pro}}+f_{t}^{\text{hp,hst,out}}-f_{t}^{\text{hp,hst,in}}+F_{mn,t}^{\text{out}}-F_{mn,t}^{\text{in}}=f_{t}^{\text{hp,sell,as}}.\hskip-5.0pt$ (33)
@@ III-B Multi-Stakeholder Sizing Model
$C_{\text{ope}}^{\text{rg}}=$ $\frac{8760}{T}\sum\nolimits_{t=1}^{T}(\sigma^{\text{deg}}P_{t}^{\text{rg,bes,d}}-P_{t}^{\text{rg,sell,hp}}\rho_{t}^{\text{rg-hp,e,rg}})\Delta t$
@@
where $\text{CRF}(r,y)={r(1+r)^{y}}/{(\left(1+r\right)^{y}-1)}$ is the capital recovery factor; $r=8\%$ and $y$ are the discount rate and lifetime; $C^{\text{rg,line}}$ is the transmission line cost; and ${\eta}_{\text{O\&M}}=2\%$ is the annual operation and maintenance (O&M) cost factor. Here, we do not plan the power lines and hydrogen pipelines; instead, their fixed costs are included in the cost function according to assumption 4).
@@
$\hskip-3.0ptC_{\text{inv}}^{\text{as}}=$ $\text{CRF}(r,y)(c^{\text{as}}W^{\text{as,asy}}+c^{\text{ast}}W^{\text{as,ast}}+c^{\text{hst}}W^{\text{as,hst}}),$ (53)
$\hskip-3.0ptC_{\text{ope}}^{\text{as}}=$ $\frac{8760}{T}\sum\nolimits_{t=1}^{T}\big[(f_{t}^{\text{as,buy,hp}}\rho_{t}^{\text{hp-as,h,as}}+P_{t}^{\text{as,buy,rg}}\rho_{t}^{\text{rg-as,e,as}})\Delta t$
@@
Given the above, we summarize the multistakeholder planning problem in Table I. For simplicity, we use $(\bm{x}^{k})_{k\in\{\text{rg,hp,as}\}}$, $(\bm{h}^{k})_{k\in\{\text{rg,hp,as}\}}$, and $(\bm{g}^{k})_{k\in\{\text{rg,hp,as}\}}$ to denote all independent decision variables, equality constraints, and inequality constraints of the RG, HP, and AS, respectively. Following this, the strategy set for each stakeholder can be defined as
$X^{k}=\{\bm{x}^{k}:\bm{h}^{k}(\bm{x}^{k})=\bm{0},\bm{g}^{k}(\bm{x}^{k})\leq\bm{0}\},~\forall k\in\{\text{rg,hp,as}\}.$ (55)
Considering the individual rationality of each stakeholder and competition in the transactions, we frame the multi-stakeholder sizing problem as a non-cooperative game and propose an MSSE model for the ReP2A system, with the game structure depicted in Fig. 2. The game is defined as follows:
  • Player: RG, HP, and AS, denoted by {rg,hp,as}.
  • Strategy: all decision variables $\bm{x}^{k}$, where $\bm{x}^{k}\in X^{k},k\in\{\text{rg,hp,as}\}$.
  • Payoff: objective function $C^{k}$, $k\in\{\text{rg,hp,as}\}$.
The competition between the RG, HP, and AS is relatively equal, which is different from the relationship between the power aggregator and electricity users [42], or between hydrogen refueling stations and hydrogen fuel cell vehicles [43]. In analogy, it is more similar to the relationship between power, nature-gas, [34, 35] and heating systems [44]. Therefore, there is no absolute leader and follower, and it is modeled as a Nash game [34, 35, 44] rather than a Stackelberg game [42, 43].
According to (46)-(54), there are conflicts of interest among RG, HP, and AS, particularly in terms of the revenues and costs related to electricity and hydrogen trading, which will be discussed in Section 10. Moreover, the investment costs of each stakeholder are not balanced due to heterogeneous physical characteristics in the staged processes, and we discuss this later in Sections V-C and V-D.
@@ IV-A Direct Approach Based on KKT Conditions
The equilibrium of noncooperative games is usually solved via fixed-point iteration [45] and direct approaches [46, 34], including those based on Karush–Kuhn–Tucker (KKT) conditions [46] and strong duality [34]. However, in the iteration method, the numerous interaction variables between stakeholders in planning problems make it difficult to ensure convergence. Additionally, the solution time for each individual stakeholder cannot be guaranteed and may even be infeasible. Thus, due to the slow convergence of fixed-point iteration, it may not be suitable for solving MSSE in this paper. Moreover, strong duality introduces bilinear terms, making it challenging to solve. Thus, we attempt to solve the MSSE model via KKT conditions.
The RG, HP, and AS models established in Section III are all convex. The KKT conditions $\Omega^{k}$ for each stakeholder $\forall k\in\{\text{rg,hp,as}\}$ are given as follows:
$\nabla_{\bm{x}^{k}}C^{k}+(\bm{\lambda}^{k})^{\text{T}}\nabla_{\bm{x}^{k}}\bm{g}^{k}+(\bm{\mu}^{k})^{\text{T}}\nabla_{\bm{x}^{k}}\bm{h}^{k}$
$+(\bm{\rho}^{\text{rg-hp,e},k})^{\text{T}}\nabla_{\bm{x}^{k}}(P^{\text{rg,sell,hp}}-P^{\text{hp,buy,rg}})$
$+(\bm{\rho}^{\text{hp-as,h},k})^{\text{T}}\nabla_{\bm{x}^{k}}(f^{\text{hp,sell,as}}-f^{\text{as,buy,hp}})=\bm{0},$ (56a)
$(\bm{\lambda}^{k})^{\text{T}}\bm{g}^{k}=\bm{0},~\bm{\lambda}^{k}\geq\bm{0},~\bm{g}^{k}\leq\bm{0},$ (56b)
$\bm{h}^{k}=\bm{0},$ (56c)
@@
where $\bm{\lambda}^{k}$ and $\bm{\mu}^{k}$ are the dual variables of $\bm{g}^{k}$ and $\bm{h}^{k}$. There are 9 dual variables in total, with 6 in (17) and 3 in (34). Among them, there are 3 for RG, 3 for HP, and 3 for AS. In summary, each KKT condition contains 3 dual variables.